OpenStudy (anonymous):

Divide as indicated. Be sure to reduce your answer to lowest terms. (Problem in comments...)

OpenStudy (anonymous):

\[\frac{ x+1 }{ x^2-16 } \div \frac{ 3x+3 }{ x+4 }\]

OpenStudy (anonymous):

@ganeshie8 , i did but i dont know how to write the final answer :/

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

@Nurali can you help my friend please ?

OpenStudy (anonymous):

pleaaaaaase lol

OpenStudy (anonymous):

\[\frac{ x+1 }{ (x+4)(x-4) } * \frac{ x+4 }{ 3(x+1) }\] x+4s cancel & x+1s cancel...I'm left with (x-4) & 3...how do I write that as my answer? 3(x-4) is incorrect & multiplying them is too.

OpenStudy (anonymous):

@whpalmer4 @campbell_st ,please help!

OpenStudy (campbell_st):

well the solution looks fine to me

OpenStudy (campbell_st):

do you need a hand with the next step...?

OpenStudy (anonymous):

do i multiply it? I thought I was done lol

OpenStudy (anonymous):

3x-12?

OpenStudy (campbell_st):